# Conformal Prediction in Hierarchical Classification

### v1

@@ 4 Experimental results
4.1 Experimental setup
@@
In this section, two types of experiments are considered in light of the concepts that have been discussed in previous sections. The first experiment demonstrates the notion of representation complexity using an example from the PlantCLEF 2015 dataset (Göeau et al., 2015). In the second experiment, different set-valued predictors are compared for solving problem (1) across four benchmark datasets, focusing on coverage, efficiency, and representation complexity. Summary statistics for these datasets are presented in Table 1. For all datasets, a predefined hierarchy $\mathcal{T}$ is extracted from the provided hierarchical labels.
All dataset images are converted to RGB format with a resolution of $200\times 200$ pixels. Hidden representations are obtained using an EfficientNet convolutional neural network (Tan & Le, 2019) pretrained on ImageNet (Krizhevsky et al., 2017). These hidden representations are fed into a hierarchical softmax layer, with dropout set to 0.1. The cross-entropy loss is minimized using the Adam optimizer, with a learning rate of $1\times 10^{-5}$ and momentum set to 0.99. Training is performed with a batch size of 32 for up to 100 epochs, with early stopping applied after five epochs of no improvement. All models are trained end-to-end on a GPU using the PyTorch library (Paszke et al., 2019), with the following hardware configuration:
@@
All inference algorithms are implemented in C++ using the PyTorch C++ API (Paszke et al., 2019).
Table 1: Summary statistics and top-1 performance for all datasets. Notation: $K$ – number of classes, $N$ – number of samples, Top-1 acc. – top-1 accuracy of hierarchical classifier.
Dataset $K$ $N_{train}$ $N_{cal}$ $N_{test}$ Top-1 acc.
CIFAR-10 10 50000 5000 5000 0.8817
Caltech-101 97 4338 2169 2169 0.9039
Caltech-256 256 14890 7445 7445 0.7578
PlantCLEF 2015 1000 91758 10723 10723 0.4156
4.2 Illustrations
(a) Cirsium arvense (L.) Scop.
(b) Lotus corniculatus L.
Figure 2: Two example images corresponding to two species from the PlantCLEF 2015 dataset. Left: Cirsium arvense (L.) Scop., right: Lotus corniculatus L..
@@ 4.2 Illustrations
Table 2: Results for CIFAR-10, Caltech-101, Caltech-256 and PlantCLEF 2015. Coverage, efficiency, and representation complexity for the following unrestricted set-valued predictors: LAC, NPS, APS, and restricted set-valued predictors: NCRSVP, CRSVP, NCRSVP-$r$, and CRSVP-$r$. The confidence level is set to 90%, and calibration and test sets are resampled 20 times. Calculated standard deviations are negligible and have therefore been excluded from the table.
Alg. LAC NPS APS NCRSVP CRSVP NCRSVP-1 CRSVP-1 NCRSVP-3 CRSVP-3
Dataset
CIFAR-10 Cov. 0.8987 0.9998 0.9008 1.0 0.9019 1.0 0.9017 0.9998 0.9016
Size 1.0412 4.9756 1.2739 7.5735 1.8418 7.7111 1.7554 5.2632 1.284
Repr. comp. 1.0357 2.5707 1.2705 1.0 1.0 1.0 1.0 1.9872 1.249
Caltech-101 Cov. 0.9025 0.9984 0.9032 0.9992 0.9012 0.9993 0.9023 0.9980 0.9036
Size 0.9322 17.114 1.5461 48.2421 5.1427 49.9799 4.9962 33.3329 2.1821
Repr. comp. 1.0 11.0477 1.5955 1.0 1.0 1.0 1.0 1.7412 1.243
Caltech-256 Cov. 0.8992 0.9821 0.9017 0.9980 0.8995 0.9993 0.9067 0.9912 0.9079
Size 2.3446 28.0667 4.7988 210.3485 43.8557 232.469 44.6294 105.4433 17.7206
Repr. comp. 2.2832 19.6563 4.4863 1.0 1.0 1.0 1.0 1.82 1.7149
PlantCLEF 2015 Cov. 0.8991 0.9183 0.8988 1.0 0.9005 1.0 0.9085 1.0 0.9134
Size 73.5715 106.05 91.8655 1000.0 676.9661 1000.0 702.7522 1000.0 591.3467
Repr. comp. 60.2026 82.6895 72.8162 1.0 1.0 1.0 1.0 1.0 1.6537
To illustrate the notion of representation complexity, we consider the PlantCLEF 2015 dataset (Göeau et al., 2015). This dataset consists of over one hundred thousand images representing 1,000 species of trees, herbs, and ferns native to the Western European region. This dataset is characterized by significant class ambiguity, making accurate predictions on species level often impossible. From an uncertainty point of view, this challenge is particularly interesting, as even biological experts sometimes encounter uncertainty when validating visual observations of living organisms.
In Figure 2, two example images are shown, corresponding to two species from the PlantCLEF 2015 dataset. The left image shows Cirsium arvense (L.) Scop., a species of the Asteraceae family, commonly known as the creeping thistle. Two set-valued predictions by means of CRSVP-$r$ (Algorithm 4) for $r=1,3$ are given as follows:
[…]
With representation complexity restricted to one, the set-valued prediction corresponds to a genus level, encompassing six species. By increasing the representation complexity to three, the set-valued prediction becomes more precise, encompassing three species of the same genus, while including the true label.
The right image shows Lotus corniculatus L., a member of the Fabaceae family, commonly referred to as bird’s-foot trefoil. Similarly, we consider two set-valued predictions obtained by CRSVP-$r$ for $r=1$ and $r=3$:
$\hat{Y}_{\leq 1}(\bm{x},u,\tau^{*})$ $=\mathcal{Y}$
[…]
In this case, the set-valued prediction with representation complexity one is imprecise, as it includes all 1,000 species. As before, increasing the representation complexity to three improves precision, i.e. narrowing the prediction set to three (visually-related) species, while including the true label.
4.3 Benchmarking results
In a second experiment, we compare various set-valued predictors that solve problem (1) across the four benchmark datasets listed in Table 1: CIFAR-10 (Krizhevsky et al., 2010), Caltech-101 (Li et al., 2003), Caltech-256 (Griffin et al., 2007), and PlantCLEF 2015 (Göeau et al., 2015). Specifically, we evaluate the following set-valued predictors: CRSVP, CRSVP-1, and CRSVP-$3$. Additionally, we demonstrate the usefulness of randomized prediction sets by considering the following naive (N) set-valued predictors: NCRSVP, NCRSVP-1, and NCRSVP-3. These correspond to setting $u$ to zero when calculating the threshold in (2.3) and constructing the prediction sets in (3.1) and (3.2), respectively. Finally, we include results for three baseline methods that produce valid set-valued predictions in flat classification (i.e. ignoring the hierarchy): the least ambiguous classifier (LAC), as proposed by Sadinle et al. (2019); adaptive prediction sets (APS), as defined in (6); and naive prediction sets (NPS), which correspond to APS without randomization (i.e. setting $u$ to zero).
The results are obtained by training a neural network with a hierarchical softmax layer, as discussed in Section 4.1. During inference, we resample the calibration and test sets 20 times and use a confidence level of 90%. The results are summarized in Table 2. For each experiment, we report coverage, efficiency, and representation complexity. Coverage is defined as the proportion of samples for which the true class is included in the prediction set. Efficiency is defined as the average size of the set-valued prediction. The results clearly indicate that naive set-valued predictors fail to deliver prediction sets with exact coverage, thereby highlighting the importance of randomized prediction sets. Moreover, increasing the representation complexity generally improves efficiency, demonstrating its practical value. In extreme cases, when representation complexity is unrestricted, such as with LAC, NPS, and APS, optimal performance in terms of efficiency is observed. However, this comes at the cost of significantly increased representation complexity in the prediction sets, in particular for large $K$, which may not be practical when predictions need to adhere to a predefined hierarchy.
5 Conclusion
In this work, we extended the split conformal prediction framework to hierarchical classification by introducing two novel set-valued prediction algorithms. The first algorithm generates valid set-valued predictions restricted to single nodes within a predefined hierarchy. We argued that this restriction can be limiting in certain applications. To address this limitation, we introduced a second algorithm that relaxes this constraint by incorporating the notion of representation complexity. Empirical evaluations on multiple benchmark datasets demonstrate the effectiveness of the proposed algorithms in achieving exact nominal coverage. In addition to exploring alternative constraints on the prediction sets, another interesting direction for future research would be to generalize our proposed methods to more complex structures, such as directed acyclic graphs.
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Supplementary materials
A.1 Computing the set of lowest common ancestors for CRSVP-$r$
In this section, we present a dynamic programming solution to (9). The pseudocode is outlined in Algorithms S5 and S6. For each internal node, $r$ local optimization problems are solved by varying the local representation complexity $r_{i}$ from 1 to $r$ and their solutions are stored. Processing children before parents ensures all necessary values are available when evaluating weak compositions within an internal node. The critical step, line 15, examines all compositions of the current $r_{i}$ into $|T|$ elements, with a runtime exponential in $r$. However, since $r$ is restricted to lower numbers, in practice, Algorithm S6 computes exact solutions efficiently.
Algorithm 5 Initialization of $s^{r}$ and $\mathcal{Q}$Input: $\bm{x},y,r,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: $s^{r}$ and $\mathcal{Q}$.
1:  $\mathcal{Q}\leftarrow\emptyset$triangleright initialize a priority queue
2:  $\mathcal{Q}\mathrm{.add}\big{(}(v_{1},\hat{P}(v_{1}\,|\,\bm{x}))\big{)}$triangleright tree root with corresponding probability mass
3:  while $\mathcal{Q}$ is not empty do
4:    $(v,\hat{p}_{v})\leftarrow\mathcal{Q}$.pop() $\triangleright$ pop the node with highest probability mass $\hat{P}(v\,|\,\bm{x})$
5:    $s^{r}(v)\leftarrow[\emptyset]\ast r$
6:    if $v$ is a leaf node then
7:       if $|\hat{\omega}(y;\bm{x})\cap v|\neq 0$ then
8:          if $\mathrm{pa}(v)\notin\mathcal{Q}$ then
9:             $\mathcal{Q}\mathrm{.add}(\mathrm{pa}(v))$
10:          end if
11:          $s^{r}(v)\leftarrow[v]\ast r$
12:       end if
13:    else
14:       for all children $v^{\prime}$ of $v$ do
15:          $\hat{p}_{v^{\prime}}\leftarrow\hat{p}_{v}\cdot\hat{P}(v^{\prime}\,|\,v,\bm{x})$triangleright compute probability mass of child node $v^{\prime}$ by means of (2)
16:          $\mathcal{Q}\mathrm{.add}\big{(}(v^{\prime},\hat{p}_{v^{\prime}})\big{)}$
17:       end for
18:    end if
19:  end while
20:  return $s^{r}$, $\mathcal{Q}$
Algorithm 6 Computation of $A_{r}(y;\bm{x})$ in (9) – Input: $\bm{x},y,r,\hat{P},\mathcal{V}_{\mathcal{T}}$
1:  $s^{r},\mathcal{Q}\leftarrow$ initialize $s^{r}$ and $\mathcal{Q}$ by means of Algorithm 5
2:  $\hat{Y}_{r}\leftarrow\emptyset$
3:  while $\mathcal{Q}$ is not empty do
4:    $v\leftarrow\mathcal{Q}$.pop()
5:    $T\leftarrow$ empty list
6:    for all children $v^{\prime}$ of $v$ do
7:       $T\leftarrow T+v^{\prime}$
8:    end for
9:    for $r_{i}=1$ to $r$ do
10:       if $|T|>r_{i}$ then
11:          $s^{r_{i}-1}(v)\leftarrow v$
12:       else
13:          $M\leftarrow$ all compositions of $r_{i}$ in $|T|$ elements
14:          $u^{*}\leftarrow+\infty$
15:          for $(i_{0},\ldots,i_{|T|-1})\in M$ do
16:             $s\leftarrow s^{i_{0}}(T[0])\cup\ldots\cup s^{i_{|T|-1}}(T[|T|-1])$
17:             if $|s|-\hat{p}_{s}<u^{*}$ then
18:               $s^{r_{i}-1}(v)\leftarrow s$
19:               $u^{*}\leftarrow|s|-\hat{p}_{s}$
20:             end if
21:          end for
22:       end if
23:    end for
24:    if $\mathrm{pa}(v)\notin\mathcal{Q}$ then
25:       $\mathcal{Q}\mathrm{.add}(\mathrm{pa}(v))$
26:    end if
27:  end while
28:  return $s^{r}(v_{1})$

### v3

@@ 4 Experimental results
Table 1: Summary statistics and top-1 performance for all datasets. Notation: $K$ – number of classes, NN – number of samples, Top-1 acc. – top-1 accuracy of classifier.
Dataset $K$ $N_{train}$ $N_{cal}$ $N_{test}$ Top-1 acc.
CIFAR-10 10 50000 5000 5000 0.8817
AMB 93 12781 1918 1917 0.8503
Caltech-101 97 4338 2169 2169 0.9039
DBpedia 219 276945 30397 30397 0.9388
Caltech-256 256 14890 7445 7445 0.7578
PlantCLEF 2015 1000 91758 10723 10723 0.4156
@@
In this section, different set-valued predictors are compared for solving problem (1) across six benchmark datasets, focusing on coverage, efficiency, and representation complexity. Summary statistics for these datasets are presented in Table 1. For all datasets, a predefined hierarchy $\mathcal{T}$ is extracted from the provided hierarchical labels. For detailed information about the experimental setup, we refer to Section A in the technical appendix.
Table 2: Results for CIFAR-10, AMB, Caltech-101, DBpedia, Caltech-256 and PlantCLEF 2015. Coverage, efficiency, and representation complexity for the following unrestricted set-valued predictors: LAC, NPS, APS, and restricted set-valued predictors: NCRSVP, CRSVP, NCRSVP-33, and CRSVP-$3$. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.
Alg. LAC NPS APS NCRSVP CRSVP NCRSVP-3 CRSVP-3
Dataset
CIFAR-10 Cov. $0.899\pm 0.005$ $0.997\pm 0.001$ $0.899\pm 0.003$ $1.000\pm 0.000$ $0.899\pm 0.005$ $0.997\pm 0.001$ $0.899\pm 0.003$
Size $1.473\pm 0.023$ $5.125\pm 0.058$ $1.849\pm 0.019$ $10.00\pm 0.000$ $3.899\pm 0.049$ $5.861\pm 0.064$ $1.946\pm 0.025$
R.C. $1.451\pm 0.021$ $3.552\pm 0.017$ $1.824\pm 0.015$ $1.000\pm 0.000$ $1.000\pm 0.000$ $2.368\pm 0.008$ $1.691\pm 0.009$
AMB Cov. $0.899\pm 0.010$ $1.000\pm 0.000$ $0.900\pm 0.009$ $1.000\pm 0.000$ $0.900\pm 0.009$ $1.000\pm 0.000$ $0.899\pm 0.009$
Size $1.128\pm 0.020$ $23.98\pm 1.195$ $1.685\pm 0.055$ $50.85\pm 1.304$ $4.856\pm 0.261$ $39.26\pm 1.574$ $2.184\pm 0.063$
R.C. $1.132\pm 0.019$ $17.38\pm 0.702$ $1.776\pm 0.051$ $1.000\pm 0.000$ $1.000\pm 0.000$ $1.829\pm 0.009$ $1.394\pm 0.018$
Caltech-101 Cov. $0.900\pm 0.007$ $1.000\pm 0.000$ $0.900\pm 0.006$ $1.000\pm 0.000$ $0.901\pm 0.005$ $1.000\pm 0.000$ $0.901\pm 0.006$
Size $0.920\pm 0.008$ $96.54\pm 0.065$ $1.165\pm 0.015$ $96.25\pm 0.049$ $4.400\pm 0.254$ $63.61\pm 0.517$ $1.784\pm 0.086$
[…]
DBpedia Cov. $0.899\pm 0.001$ $1.000\pm 0.000$ $0.901\pm 0.002$ $0.999\pm 0.000$ $0.901\pm 0.002$ $0.999\pm 0.000$ $0.901\pm 0.001$
Size $0.931\pm 0.002$ $59.17\pm 0.272$ $11.33\pm 0.287$ $162.9\pm 0.492$ $26.62\pm 0.417$ $127.9\pm 0.501$ $17.71\pm 0.423$
R.C. $1.000\pm 0.000$ $53.18\pm 0.257$ $11.90\pm 0.277$ $1.000\pm 0.000$ $1.000\pm 0.000$ $1.985\pm 0.004$ $1.380\pm 0.005$
Caltech-256 Cov. $0.900\pm 0.004$ $0.999\pm 0.000$ $0.901\pm 0.004$ $0.999\pm 0.000$ $0.900\pm 0.003$ $1.000\pm 0.000$ $0.901\pm 0.003$
Size $1.931\pm 0.040$ $62.25\pm 1.330$ $3.640\pm 0.099$ $208.4\pm 1.780$ $44.69\pm 1.252$ $163.6\pm 2.070$ $20.30\pm 0.830$
R.C. $1.926\pm 0.040$ $50.12\pm 0.996$ $3.680\pm 0.098$ $1.000\pm 0.000$ $1.000\pm 0.000$ $1.632\pm 0.005$ $1.498\pm 0.009$
PlantCLEF 2015 Cov. $0.899\pm 0.003$ $0.970\pm 0.001$ $0.899\pm 0.003$ $1.000\pm 0.000$ $0.900\pm 0.002$ $1.000\pm 0.000$ $0.899\pm 0.004$
Size $25.50\pm 0.450$ $123.6\pm 2.026$ $44.12\pm 0.994$ $998.9\pm 0.224$ $520.9\pm 4.745$ $995.3\pm 0.492$ $389.7\pm 5.898$
R.C. $24.33\pm 0.419$ $104.4\pm 1.560$ $40.19\pm 0.836$ $1.000\pm 0.000$ $1.000\pm 0.000$ $1.006\pm 0.001$ $1.632\pm 0.010$
In particular, we will compare various set-valued predictors that solve problem (1) across six datasets listed in Table 1: CIFAR-10 (Krizhevsky et al.,, 2010), Caltech-101 (Li et al.,, 2003), Caltech-256 (Griffin et al.,, 2007), PlantCLEF 2015 (Göeau et al.,, 2015), the Allen Mouse Brain (AMB) single-cell transcriptomics dataset (Tasic et al.,, 2018) and DBPedia (Ofer,, 2019). The first four datasets are image datasets from the computer vision domain. The AMB dataset contains gene expression profiles from the mouse neocortex along with corresponding cell types and a hierarchical structure. The DBpedia text dataset contains extracted structured textual content from Wikipedia. Specifically, we evaluate the following set-valued predictors: CRSVP and CRSVP-$3$. Additionally, we demonstrate the usefulness of randomized prediction sets by considering the following naïve (N) set-valued predictors: NCRSVP and NCRSVP-$3$. These correspond to setting $u$ to zero when calculating the threshold in (5) and constructing the prediction sets in (7) and (9), respectively. Finally, we include results for three baseline methods that produce valid set-valued predictions in flat classification (i.e. ignoring the hierarchy): the least ambiguous classifier (LAC), as proposed by Sadinle et al., (2019); adaptive prediction sets (APS), as defined in (6); and naïve prediction sets (NPS), which correspond to APS without randomization (i.e. setting $u$ to zero). The results presented below are obtained using a flat classifier. We do not use hierarchical classifiers, as our dynamic programming approach relies on a bottom-up traversal of the tree, and incorporating hierarchical classifiers would increase computational complexity. Moreover, Mortier et al., (2022) report no significant improvement in predictive performance when using hierarchical classifiers compared to flat ones. Nevertheless, hierarchical classifiers can still be applied within our framework, as discussed in Section 2.1. During inference, we resample the calibration and test sets 10 times and use a confidence level of 90%.
The most important results are summarized in Table 2. For each experiment, we report (marginal) coverage, efficiency, and representation complexity. Coverage is defined as the proportion of samples for which the true class is included in the prediction set. Efficiency is defined as the average size of the set-valued prediction. The results clearly indicate that naïve set-valued predictors fail to deliver prediction sets with exact coverage, thereby highlighting the importance of randomized prediction sets. Moreover, increasing the representation complexity generally improves efficiency, demonstrating its practical value. In extreme cases, when representation complexity is unrestricted, such as with LAC, NPS, and APS, optimal performance in terms of efficiency is observed. However, this comes at the cost of significantly increased representation complexity in the prediction sets, in particular for large $K$, which may not be practical when predictions need to adhere to a predefined hierarchy.
Note that imprecise predictions for low representation complexity are typically due to uncertainty in the predicted probabilities and/or characteristics of the hierarchical structure, such as inconsistency and depth. Specifically, when uncertainty spans distinct branches of the hierarchy, predictions tend to correspond to higher-level nodes for lower representation complexity. This suggests that the hierarchy may not be well-suited for certain samples. Additionally, the depth of the hierarchical structure also plays a critical role in hierarchical classification tasks with a large number of classes. In such cases, a shallow hierarchy can lead to imprecise predictions for low representation complexity, as internal nodes in shallow trees often have many children. For example, the PlantCLEF 2015 dataset, characterized by 1,000 classes and a shallow hierarchy (i.e. taxonomy consisting of a family, genus and species level), requires a higher representation complexity in order to improve efficiency.
Finally, higher representation complexity becomes valuable for datasets with many classes and high uncertainty. In this case, traditional hierarchical classification tends to return imprecise predictions, i.e. nodes higher up in the hierarchy. By restricting representation complexity, predictions can span multiple nodes lower in the hierarchy, thereby improving efficiency. This is demonstrated by comparing CRSVP and CRSVP-3 on the PlantCLEF 2015 dataset in Table 1 and Figure 3. Furthermore, Figure 3 reveals a clear trade-off between representation complexity and efficiency for the PlantCLEF 2015 dataset. In Section B of the technical appendix, we present additional results obtained for (N)CRSVP-$1$ and (N)CRSVP-$2$, and provide additional insights regarding conditional coverage and the relationship between representation complexity and efficiency.
Figure 3: Trade-off between representation complexity (log scale) and efficiency for PlantCLEF 2015. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.
5 Conclusion
In this work, we extended the split conformal prediction framework to hierarchical classification by introducing two novel set-valued prediction algorithms. The first algorithm generates valid set-valued predictions restricted to single nodes within a predefined hierarchy. We argued that this restriction can be limiting in certain applications. To address this limitation, we introduced a second algorithm that relaxes this constraint by incorporating the notion of representation complexity. Empirical evaluations on multiple benchmark datasets demonstrate the effectiveness of the proposed algorithms in achieving exact nominal coverage.
The main purpose of bounding representation complexity is to ensure the interpretability and semantic meaningfulness of predictions. However, it may also have other benefits. In particular, we conjecture that it may serve the purpose of regularization and hence increase accuracy in cases where class probabilities are poorly estimated. In such cases, flat prediction sets tend to be scattered across the entire hierarchy, and this scattering is suppressed by bounding representation complexity. We plan to elaborate on this aspect more closely in future work. Another interesting direction for future research is to generalize our methods to more complex structures, such as directed acyclic graphs.
Acknowledgements
Yusuf Sale is supported by the DAAD program Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research. Alireza Javanmardi gratefully acknowledges funding by the Klaus Tschira Stiftung (project 00.019.2024). Willem Waegeman received funding from the Flemish government under the Flanders AI research (FlAIR) program.
Supplementary Materials
Appendix A Experimental setup
We use a MobileNetV2 convolutional neural network (Sandler et al.,, 2018) pretrained on ImageNet (Krizhevsky et al.,, 2017), in order to obtain hidden representations for all image datasets. The cross-entropy loss is minimized using the Adam optimizer, with a learning rate of $1\times 10^{-5}$ and momentum set to 0.99. We set the number of epochs to 2 and 20, for the Caltech and other datasets, respectively. We train all models end-to-end on a GPU, by using the PyTorch library (Paszke et al.,, 2019) and infrastructure with the following specifications:
@@ Appendix A Experimental setup
For the AMB and DBpedia dataset, we use the same model and training procedure as described in Theunissen et al., (2024) and Mortier et al., (2021), respectively.
Appendix B Additional results
In this section, we include additional results to those presented in Table 2, namely, results for (N)CRSVP-$1$ and (N)CRSVP-$2$. In addition, we give insights into potential conditional coverage violations by examining the Pearson correlation between coverage and prediction set size. A strong correlation may indicate a potential violation of conditional coverage (Feldman et al.,, 2021). While this metric serves as a useful proxy for evaluating conditional coverage, it is not definitive, as a correlation of zero does not necessarily guarantee conditional coverage, as shown in Rossellini et al., (2024). The results are summarized in Table C3, C4 and C5.
First, note that CRSVP is different from CRSVP-$1$. The restriction is the same, however, the set of solutions for both algorithms is different: For CRSVP, starting from the mode of the distribution, the solution space is given by the mode and its ancestors. For CRSVP-$1$, the solution space is given by line 7 and 6 in Algorithm 2 and 3, respectively, namely, the most common ancestor (8) for the top-1, top-2, top-3, etc., classes. In most cases, we observe low correlation for (N)CRSVP and (N)CRSVP-$r$, reflecting strong performance regarding conditional coverage.
Finally, in Figure C4, we visually explore the relationship between representation complexity and set size for the unrestricted set-valued predictors (LAC, APS) and restricted set-valued predictors (CRSVP, CRSVP-$r$). As is clear from both Table C3, C4, C5 and Figure C4, increasing the representation complexity results in better efficiency.
@@ Appendix B Additional results
Table C3: Additional results for CIFAR-10 and AMB. Coverage, efficiency, representation complexity and Pearson correlation values for the following unrestricted set-valued predictors: LAC, NPS, APS, and restricted set-valued predictors: NCRSVP, CRSVP, NCRSVP-$r$, and CRSVP-$r$. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.
Dataset CIFAR-10 AMB
Alg. Cov. Size Repr. comp. Corr. Cov. Size Repr. comp. Corr.
LAC $0.899\pm 0.005$ $1.473\pm 0.023$ $1.451\pm 0.021$ $0.07\pm 0.011$ $0.899\pm 0.010$ $128\pm 0.020 1.1$$132\pm 0.019$ $0.060\pm 0.025$
NPS $0.997\pm 0.001$ $5.125\pm 0.058$ $3.552\pm 0.017$ $0.016\pm 0.010$ $1.000\pm 0.000$ $23.98\pm 1.195$ $17.38\pm 0.702$ -
APS $0.899\pm 0.003$ $1.849\pm 0.019$ $1.824\pm 0.015$ $0.111\pm 0.014$ $0.900\pm 0.009$ $1.685\pm 0.055$ $1.776\pm 0.051$ $0.177\pm 0.015$
NCRSVP $1.000\pm 0.000$ $10.00\pm 0.000$ $1.000\pm 0.000$ - $1.000\pm 0.000$ $50.85\pm 1.304$ $1.000\pm 0.000$ -
CRSVP $0.899\pm 0.005$ $3.899\pm 0.049$ $1.000\pm 0.000$ $0.134\pm 0.009$ $0.900\pm 0.009$ $4.856\pm 0.261$ $1.000\pm 0.000$ $0.137\pm 0.008$
NCRSVP-1 $1.000\pm 0.000$ $10.00\pm 0.000$ $1.000\pm 0.000$ - $1.000\pm 0.000$ $51.13\pm 1.338$ $1.000\pm 0.000$ -
CRSVP-1 $0.898\pm 0.005$ $3.580\pm 0.050$ $1.000\pm 0.000$ $0.207\pm 0.005$ $0.900\pm 0.008$ $4.890\pm 0.266$ $1.000\pm 0.000$ $0.148\pm 0.007$
NCRSVP-2 $0.998\pm 0.001$ $6.775\pm 0.047$ $1.616\pm 0.004$ $0.028\pm 0.009$ $1.000\pm 0.000$ $43.35\pm 1.578$ $1.511\pm 0.011$ -
CRSVP-2 $0.899\pm 0.005$ $2.279\pm 0.036$ $1.457\pm 0.005$ $0.105\pm 0.012$ $0.900\pm 0.009$ $2.666\pm 0.117$ $1.248\pm 0.011$ $0.124\pm 0.007$
NCRSVP-3 $0.997\pm 0.001$ $5.861\pm 0.064$ $2.368\pm 0.008$ $0.024\pm 0.009$ $1.000\pm 0.000$ $39.26\pm 1.574$ $1.829\pm 0.009$ -
CRSVP-3 $0.899\pm 0.003$ $1.946\pm 0.025$ $1.691\pm 0.009$ $0.103\pm 0.014$ $0.899\pm 0.009$ $2.184\pm 0.063$ $1.394\pm 0.018$ $0.114\pm 0.009$
Table C4: Additional results for Caltech-101 and DBpedia. Coverage, efficiency, representation complexity and Pearson correlation values for the following unrestricted set-valued predictors: LAC, NPS, APS, and restricted set-valued predictors: NCRSVP, CRSVP, NCRSVP-$r$, and CRSVP-$r$. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.
Dataset Caltech-101 DBpedia
Alg. Cov. Size Repr. comp. Corr. Cov. Size Repr. comp. Corr.
LAC $0.900\pm 0.007$ $0.920\pm 0.008$ $1.000\pm 0.000$ $0.886\pm 0.021$ $0.899\pm 0.001$ $0.931\pm 0.002$ $1.000\pm 0.000$ $0.814\pm 0.005$
NPS $1.000\pm 0.000$ $96.54\pm 0.065$ $1.454\pm 0.072$ - $1.000\pm 0.000$ $59.17\pm 0.272$ $53.18\pm 0.257$ $0.010\pm 0.003$
APS $0.900\pm 0.006$ $1.165\pm 0.015$ $1.251\pm 0.015$ $0.184\pm 0.013$ $0.901\pm 0.002$ $1.33\pm 0.287 11$$1.90\pm 0.277$ $0.129\pm 0.001$
NCRSVP $1.000\pm 0.000$ $96.25\pm 0.049$ $1.000\pm 0.000$ - $0.999\pm 0.000$ $162.9\pm 0.492$ $1.000\pm 0.000$ $0.019\pm 0.001$
CRSVP $0.901\pm 0.005$ $4.400\pm 0.254$ $1.000\pm 0.000$ $0.062\pm 0.009$ $0.901\pm 0.002$ $26.62\pm 0.417$ $1.000\pm 0.000$ $0.087\pm 0.002$
NCRSVP-1 $1.000\pm 0.000$ $70.36\pm 0.550$ $1.000\pm 0.000$ - $1.000\pm 0.000$ $183.2\pm 0.380$ $1.000\pm 0.000$ $0.023\pm 0.003$
CRSVP-1 $0.900\pm 0.005$ $4.173\pm 0.241$ $1.000\pm 0.000$ $0.079\pm 0.006$ $0.901\pm 0.002$ $25.59\pm 0.386$ $1.000\pm 0.000$ $0.119\pm 0.002$
NCRSVP-2 $1.000\pm 0.000$ $66.39\pm 0.528$ $1.115\pm 0.004$ - $0.999\pm 0.000$ $150.6\pm 0.418$ $1.351\pm 0.002$ $0.013\pm 0.001$
CRSVP-2 $0.900\pm 0.007$ $2.381\pm 0.110$ $1.085\pm 0.004$ $0.054\pm 0.014$ $0.901\pm 0.002$ $9.34\pm 0.484 1.1$$191\pm 0.002$ $0.107\pm 0.001$
NCRSVP-3 $1.000\pm 0.000$ $63.61\pm 0.517$ $1.306\pm 0.008$ - $0.999\pm 0.000$ $127.9\pm 0.501$ $1.985\pm 0.004$ $0.004\pm 0.001$
CRSVP-3 $0.901\pm 0.006$ $784\pm 0.086 1.1$$133\pm 0.007$ $0.066\pm 0.011$ $0.901\pm 0.001$ $17.71\pm 0.423$ $1.380\pm 0.005$ $0.113\pm 0.001$
Table C5: Additional results for Caltech-256 and PlantCLEF 2015. Coverage, efficiency, representation complexity and Pearson correlation values for the following unrestricted set-valued predictors: LAC, NPS, APS, and restricted set-valued predictors: NCRSVP, CRSVP, NCRSVP-$r$, and CRSVP-$r$. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.
Dataset Caltech-256 PlantCLEF 2015
Alg. Cov. Size Repr. comp. Corr. Cov. Size Repr. comp. Corr.
LAC $0.900\pm 0.004$ $1.931\pm 0.040$ $1.926\pm 0.040$ $0.29\pm 0.006$ $0.899\pm 0.003$ $25.50\pm 0.450$ $24.33\pm 0.419$ $0.20\pm 0.009$
NPS $0.999\pm 0.000$ $62.25\pm 1.330$ $50.12\pm 0.996$ $0.01\pm 0.009$ $0.970\pm 0.001$ $123.6\pm 2.026$ $104.4\pm 1.560$ $0.003\pm 0.006$
APS $0.901\pm 0.004$ $3.640\pm 0.099$ $3.680\pm 0.098$ $0.05\pm 0.008$ $0.899\pm 0.003$ $44.12\pm 0.994$ $40.19\pm 0.836$ $0.01\pm 0.010$
NCRSVP $0.999\pm 0.000$ $208.4\pm 1.780$ $1.000\pm 0.000$ - $1.000\pm 0.000$ $998.9\pm 0.224$ $1.000\pm 0.000$ -
CRSVP $0.900\pm 0.003$ $44.69\pm 1.252$ $1.000\pm 0.000$ $0.100\pm 0.007$ $0.900\pm 0.002$ $520.9\pm 4.745$ $1.000\pm 0.000$ $0.340\pm 0.004$
NCRSVP-1 $0.999\pm 0.000$ $215.5\pm 1.885$ $1.000\pm 0.000$ - $1.000\pm 0.000$ $998.9\pm 0.176$ $1.000\pm 0.000$ -
CRSVP-1 $0.901\pm 0.003$ $44.24\pm 1.297$ $1.000\pm 0.000$ $0.139\pm 0.004$ $0.900\pm 0.003$ $534.5\pm 4.492$ $1.000\pm 0.000$ $0.356\pm 0.005$
NCRSVP-2 $0.999\pm 0.000$ $179.0\pm 2.035$ $1.242\pm 0.004$ - $1.000\pm 0.000$ $997.0\pm 0.443$ $1.002\pm 0.001$ -
CRSVP-2 $0.901\pm 0.003$ $27.63\pm 0.898$ $1.270\pm 0.006$ $0.073\pm 0.007$ $0.898\pm 0.003$ $442.1\pm 4.337$ $1.290\pm 0.004$ $0.295\pm 0.004$
NCRSVP-3 $1.000\pm 0.000$ $163.6\pm 2.070$ $1.632\pm 0.005$ - $1.000\pm 0.000$ $995.3\pm 0.492$ $1.006\pm 0.001$ -
CRSVP-3 $0.901\pm 0.003$ $20.30\pm 0.830$ $1.498\pm 0.009$ $0.045\pm 0.009$ $0.899\pm 0.004$ $389.7\pm 5.898$ $1.632\pm 0.010$ $0.262\pm 0.004$
Figure C4: Representation complexity versus set size for the following unrestricted set-valued predictors: LAC and APS, and restricted set-valued predictors: CRSVP and CRSVP-$r$, for all datasets. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.